OpenStudy (mellamokatie):

Each of two urns contains green balls and red balls. Urn I contains 8 green balls and 12 red balls. Urn II contains 5 green balls and 8 red balls. If a ball is drawn from each urn, what is P(red and red)? A. 79/65 B. 24/65 C. 20/30 D. 2/13

OpenStudy (parriscite):

U_1 = 12/20 || U_2 = 8/13 The probability of two events happening is their product. After you simplify the fraction, you will find the answer.

OpenStudy (kropot72):

The events 'draw a ball from urn I' and 'draw a ball from urn II' are independent. Therefore the probability P(red and red) is the product of the two probabilities: \[\large P(red\ with\ red)=\frac{12}{20}\times\frac{8}{13}=? \]

OpenStudy (mellamokatie):

one moment

OpenStudy (mellamokatie):

.3?

OpenStudy (mellamokatie):

@Mehek14

OpenStudy (mellamokatie):

HI! Thank god someone is here haha

563blackghost (563blackghost):

Heyo! What @kropot72 is quite right and your answer is sort of correct....it would actually be `3.692307...` in which it continues on. Which answer have you chosen?

OpenStudy (mellamokatie):

Yeah I dont know honestly lol

563blackghost (563blackghost):

Well we would first multiply the fractions of course... \(\Huge{\frac{12}{20} \times \frac{8}{13} = \frac{96}{260}}\) Now we would simplify this....what number would be simplify by? Factors of 260: 1,2,4,5,10,13,20,26,52,65,130,260 Factors of 96: 1,2,3,4,6,8,12,16,24,32,48,96 The factor we will use is the one they both have in common.

OpenStudy (mellamokatie):

So t would be 2?

563blackghost (563blackghost):

Not quite. It would actually be 4 :) So we divide by 4... \(\Huge{\frac{96}{260} \div \frac{4}{4} = ?}\)

OpenStudy (mellamokatie):

ohhh thats right the last one

OpenStudy (mellamokatie):

1

563blackghost (563blackghost):

Last one? Do you mean D?

OpenStudy (mellamokatie):

No the last number as in 4

563blackghost (563blackghost):

Ooo ok

OpenStudy (mellamokatie):

So what would it be?

563blackghost (563blackghost):

Well we would divide by 4... \(\Huge{\frac{96 \div 4}{260 \div 4}=?}\) What did you get?

OpenStudy (mellamokatie):

24/ 65

563blackghost (563blackghost):

So the answer is B ^^

OpenStudy (mellamokatie):

Oh awesome thank you can you answer some more?